OpenStudy (anonymous):

Michael rolled two number cubes, numbered 1 to 6, simultaneously. According to theoretical probability, how many of the rolls should Michael expect the sum to be 6 if he rolled them 108 times?

OpenStudy (anonymous):

sorry i suck at probability!

OpenStudy (anonymous):

me too:(

OpenStudy (anonymous):

do you think i would divide 6/108?

OpenStudy (anonymous):

i think u find all the possibilites that would make it 6, and then out of 36 (since that's the total number of sums.

OpenStudy (anonymous):

n=36 so do we have to get the score on the upperface of both the cube be 6

OpenStudy (anonymous):

1+5,2+4,3+3,4+2,5+1 (these are 5 possibilites) (5/36)^108 ---> i think :D

OpenStudy (zarkon):

\[\frac{5}{36}\cdot 108=15\]

OpenStudy (anonymous):

@zarkon is perfect

OpenStudy (anonymous):

thank you!!!

OpenStudy (anonymous):

yeah i meant to multiplication sign. sorry! yeah that means i got it right too!

OpenStudy (anonymous):

welcome

OpenStudy (anonymous):

haha i gave all of you guys medals!

OpenStudy (anonymous):

now i will give u a medal

OpenStudy (zarkon):

this is binomial...the mean is \(n\cdot p\)

OpenStudy (anonymous):

zarkon's in the lead for medals :D

OpenStudy (anonymous):

Yeahhh!!!! haha medals for evryone:D